The triangle is 48 square inches and the height is 12 Inches. How long is the base of the triangle
Answer by checkley79(3341)

AREA=BH/2
48=B*12/2
40=6B
B=40/6
B=20/3=6.67 INCHES FOR THE BASE.
PROOF:
48=6.67*12/2
48=80/2
40=40